OpenStudy (anonymous):

Please answer this for me!! y = log x If y = 10, then what is x?

OpenStudy (anonymous):

\[\log_{x}a = b \rightarrow a = x ^{b} \]

OpenStudy (anonymous):

so, \[\log_{10}x = y \rightarrow x =? \]

OpenStudy (anonymous):

Raise 10 to the power of both sides of the equation.

OpenStudy (anonymous):

is this right? y = log x 10 = log x 10 = 10^x x = 1

OpenStudy (anonymous):

10 = log x so x = ? follow the logic of what I posted earlier.

OpenStudy (anonymous):

10*10

OpenStudy (anonymous):

your answer is incorrect ,BTW

OpenStudy (anonymous):

damn

OpenStudy (anonymous):

x=10^10.............any dobt in this...!!!

OpenStudy (anonymous):

\[\log _{a}X=Y where X=a ^{Y}\] First guy wrote it the right way but he used X and Y in a different way which can confuse you. Also if not told in the question. a =10 and you know Y= 10 so just plug it in.

OpenStudy (anonymous):

thanks i think i get it now